1)If a maximum interest rate for credit cards is set above the equilibrium interest rate, ________________.

interest rates will increase

interest rates will decrease

there will be a surplus of credit cards

2)Janush is leaving his home country of Zendra to study for his economics degree in another country. He plans to return to his country when he finishes his studies in 4 years.  If Zendra's present GDP is $835,730 and grows 3% every year, what is the future GDP of Zendra 4 years from now? (round to the nearest dollar)

3)Convergence is the idea that ______________________.

high-income countries grow faster than low-income countries

low-income countries grow faster than high-income countries

low-income countries grow at the same rate as high-income countries

none of the above.

4)Suppose political instability in Russia depreciates its currency to a low of 68.2 RUB=$1. At the same time, this instability causes the Russian economy to enter a recession and GDP falls from 190,000 RUB to 140,000 RUB. What is the current value of Russia's GDP measured in USD? Round your answer to the nearest hundredth.
Answer #1

1.

Correct option: there will be a surplus of credit cards

Reason: When interest rate on credit cards is higher than equilibrium interest rates, then supply of credit cards would exceed demand, leading to surplus of credit cards.

2.

FV of Zendra GDP after 4 year = GDP(1+r)n

Substitute given values,

GDP after 4 years = 835,730(1+0.03)4

GDP after 4 years = 940621.5

3.

Correct option: low-income countries grow faster than high-income countries

Reason: The idea of convergence states that low-income countries grow faster than high-income countries, thereby converging into high income countries growth rates

4.

Since 1$ = 68.2 RUB,

Wen GDP = 140,000 RUB

Then this can be expressed in dollars as:

GDP = 140,000/68.2

GDP = $2053 approximately